### v3

# Fluctuation Theorem and Optimal Control of an Active Tracking Particle with Information Processing

Abstract
Living systems often function with regulatory interactions, but the question of how activity, stochasticity and regulations work together for achieving different goals still remains puzzling. We propose the model of an active tracking particle with information processing, based on which the entropy production, information flow, and generalised fluctuation theorem are derived. Moreover, the system performance, in terms of the first passage steps and the total energy consumption, are analysed in the variable space of (measurement error, control field), leading to discussions on the optimal control of the system. Not only elucidating the basic concepts involved in a stochastic active system with information processing, this prototypical model could also inspire more elaborated modelings of natural smart organisms and industrial designs of controllable active systems with desired physical performances in the future.
Among the theoretical concepts proposed for understanding physics of life, active matter can be regarded as a very successful trial [1, 2, 3, 4]. Not only now serving as a model system for studying the far-from-equilibrium statistical physics, active matter, which is often assumed with simple physical rules, also keeps refreshing our understandings of the living systems, in terms of both the individual behaviours such as microorganism locomotion [5, 6, 7, 8] and the collective responses such as bird flocking [9, 10, 11, 12]. Although different active models have gained great successes in analysing the physical responses in customised cases, such models are usually oversimplified for describing the biological activities of most creatures with complicated regulatory interactions, which often include signal sensing, decision making and adaptive responses [13, 14, 15, 16].
In living systems, the acquired knowledge of their own internal status or the external environment can be utilised to modulate their responses, i.e., regulations concerning information flow, and such regulatory examples range from the microscopic ones such as algae phototaxis [17, 18] to the macroscopic ones such as human crowding [19, 20]. Inspired by such examples, there have been various studies which implement external controls to achieve the desired physical performances of soft or active matter systems [21, 22, 23, 24], e.g., passive-to-active transformation of a Brownian particle [25, 26], microswimmer navigation with minimum time or energy consumption [27, 28, 29, 30, 31, 32, 33, 34], and maximum work efficiency of information engines [35, 36, 37]. However, despite these initial attempts illustrating the importance of regulations in controlling soft or active matter systems, how to formulate the information flow, entropy production and physical performances in a unified theoretical framework still remains questionable, rendering many difficulties in understanding the responses of a regulated living system from a physical perspective.
Figure 1: (a) Evolution of the system consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation, and (b) zoom-in of the system dynamics during the time step, $t_{k}\rightarrow t_{k+1}$.
In this work, we propose the model of an active tracking particle with information processing, based on a Bayesian description. With the model, we obtain the generalised fluctuation theorem, serving as the basic obeying principle of the information flow and entropy production. Moreover, we investigate how the performance of the system, in terms of the first passage steps and the energy consumptions, can depend on the information-related processes including measurements and feedback controls, providing the optimal control protocols in the corresponding parameter space.
System setup.
For seeking the transparency of the model, we take a one-dimensional and descritised description to illustrate how a regulated active particle should move, as shown in Fig. 1. The state of the active particle at any time, $t_{k\in\mathbb{Z}}$, is described by two variables $(x_{k},n_{k})$, where $x_{k}=s~\!\Delta x$ denotes its location with $s\in\mathbb{Z}$ and $\Delta x$ as the step length, and $n_{k}\in(L,R)$ denotes its orientation pointing to the left or the right; without the information processing, the model reduces to the run-and-tumble model describing the dynamics of a bacterium [38, 39]. The regulatory purpose of the active particle is to move from the location $x=0$ to the destination $x=\mathcal{D}\Delta x$ along a straight line, i.e., the location should be $x=d~\!\Delta x$ at any time $t_{d}$, and this is a typical tracking problem in control theories. To achieve feedback controls, we consider the particle carrying a magnetic moment of magnitude $m$ in the same direction as its orientation, and the orientation can be controlled by applying an external magnetic field; such a model can be easily generalised to other controlling cases, e.g., by utilising light signal on algae [17, 18] or chemical concentration on bacteria [40, 41]. As shown in Fig. 1, the particle starts moving from state $(x_{0},n_{0})$ at time $t_{0}$, and after measurement - feedback control- active motion - relaxation, the particle reaches a new state $(x_{1},n_{1})$ at time $t_{1}$. By repeating the above process, the particle stops moving after arriving at the destination $(\mathcal{D}\Delta x,n_{\mathcal{T}})$ after $\mathcal{T}$ time steps ($\mathcal{T}\geq\mathcal{D}$); note that the particle motion can deviate from the desired trajectory (straight line) due to the intrinsic stochasticity of the system.
We utilise a bipartite Bayesian network [42, 43] to describe the dynamics of the active particle, consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion as an ordinary run-and-tumble particle and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation. With this treatment, the regulation of the whole process simply follows the classical Bellman description [44] of dynamic programming where the total optimal cost is the summation of all the costs at every time step, so it is sufficient to focus on the regulation process during a single time step, $t_{k}\rightarrow t_{k+1}$ with $0\leq k<\mathcal{T}$. For achieving regulatory purposes during any arbitrary time step, $t_{k}\rightarrow t_{k+1}$, we take the Maxwell-demon-like setup of the measurement - feedback process [45, 46, 47], and it is detailed as follows. At time $t_{k}$, the state of the particle is $(x_{k},n_{k})$, where $n_{k}=L,R$ with equal probability at equilibrium, i.e., $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=1/2$. Then we measure its orientation denoted as $m_{k}$, and such measurements have an error, $0\leq\epsilon\leq 1$; in other words, we can define a conditional probability characterising the measurement: $p(m_{k}|n_{k})=1-\epsilon$ for $m_{k}=n_{k}$ and $p(m_{k}|n_{k})=\epsilon$ for $m_{k}\neq n_{k}$. If the measured orientation is pointing to the left, i.e., $m_{k}=L$, then we apply a magnetic field of the magnitude $B$ and the direction pointing to the right during a finite time interval $\Delta t$ to change its orientation; if the measured orientation is pointing to the right, i.e., $m_{k}=R$, then we do not apply the magnetic field and let the particle orientation freely evolve during $\Delta t$ by thermal fluctuation. We use $n_{k}^{\prime}$ to denote the particle orientation after the measurement and feedback control. The transition, $n_{k}\rightarrow n_{k}^{\prime}$, which is critical in the regulatory process, includes the nodes, $n_{k}$, $m_{k}$, and $n_{k}^{\prime}$ as shown in Fig. 1, and the transition can be described by the conditional probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, with its explicit expressions shown in the table of the Appendix. For example, no magnetic field is applied in the case of ($n_{k}=L$, $m_{k}=R$), where the particle orientation will evolve freely during $\Delta t$, and we can obtain the probabilities of finding the particle with $n_{k}^{\prime}=L$ and $n_{k}^{\prime}=R$ as, $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$ and $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$ without magnetic field, respectively. In following discussions, the notation $u_{p}=\exp[-k_{0}\Delta t]$ (related with unchanged population) is taken for simplicity. After the measurement - feedback process, we let the particle move along its regulated orientation with a constant step length, $\Delta x$, which is described by the process $(x_{k},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k}^{\prime})$, and such dynamic process can be achieved in various active systems, e.g., by shining light on photocatalytical colloidal particles [48, 49] or applying electric field on metal-dielectric Janus particles [50, 51]. Then let the particle orientation relax to the state with its orientation distribution recovering to $p_{\mathrm{eq}}^{0}(n_{k+1}=L,R)=1/2$ at time $t_{k+1}$. For those familiar with bacteria swimming in the run-and-tumble mode, the above setup can be treated as a bacterium moving with a ‘brain’ which performs measurement and feedback control. By repeating the above process from $t_{0}$ to $t_{\mathcal{T}}$, the particle can arrive at its destination after $\mathcal{T}$ time steps, with its state as $(\mathcal{D}\Delta x,n_{\mathcal{T}})$.
Entropy production, information flow and fluctuation theorem.
The total entropy production in the whole system consists of the entropy production in both the subsystem $\mathcal{X}$ and its connected heat bath [42, 52], i.e., $\Delta\mathcal{S}_{\mathrm{tot}}=\Delta\mathcal{S}_{\mathcal{X}}+\Delta\mathcal{S}_{\mathrm{bath}}=\sum_{k=0}^{\mathcal{T}-1}(\Delta s_{\mathcal{X}}^{k}+\Delta s_{\mathrm{bath}}^{k})$, where $\Delta s_{\mathcal{X}}^{k}$ and $\Delta s_{\mathrm{bath}}^{k}$ denote the entropy production of the subsystem $\mathcal{X}$ and the heat bath during the time step, $t_{k}\rightarrow t_{k+1}$, respectively. During any time step, $t_{k}\rightarrow t_{k+1}$, the entropy production of the subsystem $\mathcal{X}$ is simply,
$\Delta s_{\mathcal{X}}^{k}=\ln\frac{p(n_{k})}{p(n_{k+1})},$ (1)
where $p(n_{k})$ denotes the probability distribution function of the particle orientation at time $t_{k}$. For explicit expressions of relevant probabilities in the system, one can refer to supplemental materials [53]. Regarding the entropy change of the heat bath, $\Delta s_{\mathrm{bath}}^{k}$, during the time step, $t_{k}\rightarrow t_{k+1}$, which is due to the heat dissipated into the heat bath from the system $\mathcal{X}$, there are two parts: one during the process $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, and the other during the process $(x_{k+1},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k+1})$, which are denoted by $\Delta s_{\mathrm{bath}}^{k1}$ and $\Delta s_{\mathrm{bath}}^{k2}$, respectively. The entropy production, $\Delta s_{\mathrm{bath}}^{k1}$, during the process, $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, can be calculated as:
$\Delta s_{\mathrm{bath}}^{k1}=-\beta Q_{k}^{1}=\ln\frac{p(n^{\prime}_{k}|n_{k},m_{k})}{p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})},$ (2)
where $\beta=1/k_{B}T$ with $T$ as the temperature, $Q_{k}^{1}$ denotes the dissipated heat during the process $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, and $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$ denotes the backward transition probability from $\hat{n}^{\prime}_{k}$ to $\hat{n}_{k}$ depending on the memory $\hat{m}_{k}$, with $\hat{n}_{k}$, $\hat{n}^{\prime}_{k}$ and $\hat{m}_{k}$ as time-reversed $n_{k}$, $n^{\prime}_{k}$ and $m_{k}$, respectively. The backward transition probability, $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$, can be related with the forward transition probability, $p(n^{\prime}_{k}|n_{k};m_{k})$, generally as: $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})/p(n^{\prime}_{k}|n_{k};\lambda_{k})=\exp[-\beta E(n_{k};\lambda_{k})+\beta E(n^{\prime}_{k};\lambda_{k})]$, where $\lambda_{k}$ is external control parameter characterising the control on $\mathcal{X}$ and $E(n_{k};\lambda_{k})$ denotes the system energy. Without magnetic field, we have degenerated states: $E(n_{k};\lambda_{k})=E(n^{\prime}_{k};\lambda_{k})=0$, and there are relations, $\exp[-\beta E(n_{k};\lambda_{k})]=\exp[-\beta E(n^{\prime}_{k};\lambda_{k})]=1$ and $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})=p(n^{\prime}_{k}|n_{k};\lambda_{k})$. With magnetic field, the degeneracy is lifted, and we have $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})\neq p(n^{\prime}_{k}|n_{k};\lambda_{k})$, whose values depend on $\hat{n}_{k}$ and $\hat{n}^{\prime}_{k}$. The entropy production, $\Delta s_{\mathrm{bath}}^{k2}$, during another process, $(x_{k+1},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k+1})$, is simply:
$\Delta s_{\mathrm{bath}}^{k2}=-\beta Q_{k}^{2}=\ln\frac{p(n_{k+1}|n^{\prime}_{k})}{p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})}.$ (3)
Then the total entropy production during the time step, $t_{k}\rightarrow t_{k+1}$, becomes:
$\sigma^{k}$ $=$ $\Delta s_{\mathcal{X}}^{k}+\Delta s_{\mathrm{bath}}^{k1}+\Delta s_{\mathrm{bath}}^{k2}$ (4)
$=$ $\ln\frac{p(n_{k})p(n^{\prime}_{k}|n_{k},m_{k})p(n_{k+1}|n^{\prime}_{k})}{p(n_{k+1})p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})}.$
One can easily prove that, if there is no measurement and feedback control, the total entropy production satisfies the integral fluctuation theorem: $\langle e^{-\Delta\mathcal{S}_{\mathrm{tot}}}\rangle=\langle e^{-\sum_{k=0}^{\mathcal{T}-1}\sigma^{k}}\rangle=\prod_{k=0}^{\mathcal{T}-1}\left[\sum_{\hat{n}_{k},\hat{n}^{\prime}_{k},n_{k+1}}p(n_{k+1})p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})\right]=1$, where $\langle$$\rangle$ indicates the ensemble average and the last equality is guaranteed by the normalisation condition. Then there is the second law of thermodynamics: $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq 0$ obtained from Jensen’s inequality. By considering the measurement and feedback control, the relation $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq 0$ may not hold any more. For characterising how measurement and feedback control can influence the dynamics of the active particle, we introduce an information quantity, $\Theta_{k}$ [42]:
$\Theta_{k}=I_{k}^{\mathrm{fin}}-I_{k}^{\mathrm{tr}}-I_{k}^{\mathrm{ini}},$ (5)
where the mutual information $I_{k}^{\mathrm{fin}}=I(n^{\prime}_{k}:m_{k})=\ln[p(n^{\prime}_{k},m_{k})/p(n^{\prime}_{k})p(m_{k})]$ characterises the correlation between $n^{\prime}_{k}$ and $m_{k}$, the mutual information $I_{k}^{\mathrm{tr}}=I(m_{k}:n_{k})=\ln[p(m_{k},n_{k})/p(m_{k})p(n_{k})]$ denotes the transfer entropy from the physical system $\mathcal{X}$ to the control system $\mathcal{C}$, and $I_{k}^{\mathrm{ini}}=0$ denotes no correlation between $n_{k}$ and previous measurements, $m_{l<k}$. By summing them up, we obtain the dynamic information flow in the system, $\Theta_{d}=\sum_{k=0}^{\mathcal{T}-1}\Theta_{k}$. One can prove that the total entropy production, $\Delta\mathcal{S}_{\mathrm{tot}}$, and the dynamic information flow, $\Theta_{d}$, satisfy the generalised integral fluctuation theorem:
$\langle e^{-\Delta\mathcal{S}_{\mathrm{tot}}+\Theta_{d}}\rangle=$ (6)
$~\!\!\prod_{k=0}^{\mathcal{T}-1}\Big{[}\!\!\!\sum_{n_{k},m_{k}}^{n^{\prime}_{k},n_{k+1}}\frac{p_{B}(\hat{n}_{k+1},\hat{n}^{\prime}_{k},\hat{m}_{k},\hat{n}_{k})}{p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})}p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})\!\Big{]}\!=\!1,$
as obtained in other information-incorporated stochastic systems [54, 55, 56, 42], where the forward probability is $p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})\quad=\quad p(n_{k})p(m_{k}|n_{k})p(n^{\prime}_{k}|n_{k},m_{k})\times\\ p(n_{k+1}|n^{\prime}_{k})$ and the backward probability is $p_{B}(\hat{n}_{k},\hat{m}_{k},\hat{n}^{\prime}_{k},\hat{n}_{k+1})=p(n_{k+1})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})p(m_{k}|n^{\prime}_{k})\times\\ p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$. Alternatively, one can utilise the Kullback-Leibler (KL) divergence, $d_{\mathrm{KL}}=\Delta\mathcal{S}_{\mathrm{tot}}-\Theta_{d}=\ln[p(\mathcal{X}\bigcup\mathcal{C})/p_{B}(\mathcal{X}\bigcup\mathcal{C})]$, to express the generalised fluctuation theorem: $\langle e^{-d_{\mathrm{KL}}}\rangle$=1. By again applying Jensen’s inequality, we can obtain the generalised second law of thermodynamics, $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq\langle\Theta_{d}\rangle$ [42, 57], indicating that the total entropy production $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle$ can be negative.
Figure 2: Mean first passage steps $\langle\mathcal{T}\rangle$ as a function of the magnetic field $\hat{B}$, for $\epsilon=0$ (red), 1/2 (blue) and 1 (brown). The black solid line is obtained by $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$ for $\hat{B}\rightarrow 0$, and the round circles for different $\epsilon$ denote the values obtained by $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$ for $\hat{B}\rightarrow\infty$. The parameter, $u_{p}=\exp[-k_{0}\Delta t]=0.9$ is taken for illustration.
Figure 3: Mean energy consumption per time step, $\ln[\langle\mathcal{E}\rangle/(\mathcal{D}E_{A}^{0})]$, as a function of the measurement error and the control field for parameters of (a) $\mathcal{B}=1/2,\mathcal{M}=1$, (b) $\mathcal{B}=1/2$, $\mathcal{M}=30$, and (c) $\mathcal{B}=10$, $\mathcal{M}=0.1$, where the black circles denote the local minima and the red triangle denote the global minimum. (d) Optimal measurement error and (e) optimal field as a function of the measurement parameter, $\mathcal{M}$, for the control parameter of $\mathcal{B}=1/2$ (black line), 1010 (red line) and 5050 (blue line). Squares and circles in (e) denote the values of the optimal field at the limit of $\mathcal{M}\rightarrow 0$ and $\mathcal{M}\rightarrow\infty$, respectively. (f) Optimal field at the limit of $\mathcal{M}\rightarrow 0~(\epsilon=0)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}$, and (g) at the limit of $\mathcal{M}\rightarrow\infty~(\epsilon=1)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}$, as a function of the control parameter, $\mathcal{B}$. Inset of (f) denote the values of the optimal field for the full range of $\mathcal{B}\in(10^{-6},5\times 10^{1})$.
First passage steps – $\mathcal{T}$.
Due to the stochastic nature of the system, the active particle moves along different trajectories to its destination with different first passage steps, $\mathcal{T}$, minimising the average of which can be an important regulation goal. Suppose that after $\mathcal{T}$ steps, the particle reaches the destination $x=\mathcal{D}\Delta x$. Then the particle must move to the right with $\mathcal{W}+\mathcal{D}$ steps and to the left with $\mathcal{W}$ steps, where $\mathcal{T}=2\mathcal{W}+\mathcal{D}$. Here, the total number of possible trajectories to reach $\mathcal{D}$ for the first time after $\mathcal{T}$ steps is $\mathcal{D}\cdot C_{\mathcal{T}}^{\mathcal{W}}/\mathcal{T}$, resembling the result of Catalan’s trapezoid problem [58]. Correspondingly, we can obtain the probability distribution of the first passage steps as:
$P(\mathcal{T})=\frac{\mathcal{D}}{\mathcal{T}}C_{\mathcal{T}}^{\mathcal{W}}[p(n^{\prime}_{k}=R)]^{\mathcal{W}+\mathcal{D}}[p(n^{\prime}_{k}=L)]^{\mathcal{W}},$ (7)
where $p(n^{\prime}_{k}=R)$ and $p(n^{\prime}_{k}=L)=1-p(n^{\prime}_{k}=R)$ denotes the probability of the particle moving to the right and the left, respectively, and it satisfies the normalisation condition: $\sum_{\mathcal{T}=\mathcal{D}}^{\infty}P(\mathcal{T})=1$.
The average of the first passage steps is
$\langle\mathcal{T}\rangle$ $=$ $\sum_{\mathcal{T}=\mathcal{D}}^{\infty}\mathcal{T}\cdot\frac{\mathcal{D}}{\mathcal{T}}C_{\mathcal{T}}^{\mathcal{W}}[p(n^{\prime}_{k}=R)]^{\mathcal{W}+\mathcal{D}}[p(n^{\prime}_{k}=L)]^{\mathcal{W}}$
$=$ $\frac{\mathcal{D}}{2p(n^{\prime}_{k}=R)-1},$ (8)
which depends on the process of measurement and feedback control through the probability, $p(n^{\prime}_{k}=L)$ [explicit expression shown in the table of the Appendix]. Note that $\langle\mathcal{T}\rangle$ diverges for unbiased random walk with $p(n^{\prime}_{k}=R)=1/2$. Apparently, the averaged first passage steps decreases with increasing magnetic field and decreasing measurement error, as shown in Fig. 2. We introduce the dimensionless field strength, $\hat{B}=B/B_{0}$ with $B_{0}=k_{B}T/m$ as the referenced magnetic field. For weak field controls, $\hat{B}\rightarrow 0$, the averaged first passage steps decreases with the increasing magnetic field in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$, and for strong field controls, $\hat{B}\rightarrow\infty$, the averaged first passage steps only depends on the measurement error, in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$.
Energy consumption – $\mathcal{E}$.
Suppose that during each time step, $t_{k}\rightarrow t_{k+1}$, the energy consumed for sustaining the particle activity is $E_{A}^{0}$, which is usually proportional to $v_{0}^{2}$ with $v_{0}$ as self-propulsion speed, the energy of applied magnetic field is $E_{B}=E_{B}^{0}\cdot(B/B_{0})^{2}=E_{B}^{0}\hat{B}^{2}$, and the energy for measurement is $E_{M}=-E_{M}^{0}\ln\epsilon$, which increases with the decreasing error. Then the ensemble-averaged total energy consumed for arriving at the destination is,
$\langle\mathcal{E}\rangle$ $=$ $\frac{\mathcal{D}E_{A}^{0}}{2p(n^{\prime}_{k}=R)-1}~\left[1+\mathcal{B}\hat{B}^{2}\cdot p(m_{k}=L)-\mathcal{M}\ln\epsilon\right],$
where $p(m_{k}=L)=1/2$ denotes the probability of the measured particle orientation as $m_{k}=L$, and the referenced magnetic energy, $\mathcal{B}=E_{B}^{0}/E_{A}^{0}$, and referenced measurement energy, $\mathcal{M}=E_{M}^{0}/E_{A}^{0}$, are two system parameters.
Note that the measurement energy, $E_{M}$, increases with the decreasing error in measurements, and meanwhile, the decreasing error can shorten the total time through biasing the probability $p(n^{\prime}_{k}=R)$ and thus lowers the energy spent on the active motion and the control field; such a competition can lead to the issue of (information) robustness - (measurement) energy tradeoff, commonly discussed in biological systems [59, 60, 61]. As shown in Fig. 3(a), for system parameters of $\mathcal{B}=1/2$ and $\mathcal{M}=1$, the total consumed energy can have a local minimum (also the global minimum in this case) at $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$, which serves as the optimal control variables for minimizing the total energy consumption. By increasing the referenced magnetic energy, e.g., $\mathcal{M}=30$ in Fig. 3(b), the local minimum is no longer the global one, where the global minimum is now located at $\epsilon=1$ which is a natural upper bound of the measurement error. Note that the global minimum located at $\epsilon=1$ is not necessarily an extremum (usually not). Meanwhile, the situation can become more complicated at large $\mathcal{B}$, where there can be two local minima, as shown in Fig. 3(c). Then the global minimum is chosen among the local minima and the one at $\epsilon=1$, which is shown in Fig. 3(d,e). In this case, the system can exhibit a transition of controllable variables $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$ depending on the system parameters $(\mathcal{B},\mathcal{M})$, where the guiding rule is to minimize the total energy consumption. There are two important limits to discuss about, $\mathcal{M}=0$ and $\mathcal{M}\rightarrow\infty$. For the measurement parameter, $\mathcal{M}=0$, the measurements of optimal controls are error-free, i.e., $\epsilon^{\mathrm{opt}}=0$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\frac{1}{2}W(4/\mathcal{B})$ [$W(x)$ as the principal branch of Lambert W function] at $\mathcal{B}\rightarrow 0$ [blue line in Fig. 3(f)], and $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$ [red line in Fig. 3(f)]. For the measurement parameter, $\mathcal{M}\rightarrow\infty$, the measurements of optimal controls are wrong, i.e., $\epsilon^{\mathrm{opt}}=1$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\frac{1}{2}W(8/[\mathcal{B}\mathcal{(}1-u_{p})])$ at $\mathcal{B}\rightarrow 0$ [blue line in Fig. 3(g)], and $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$ [red line in Fig. 3(g)].
Summary.
In this work, we develop a smart active model where activity, thermal fluctuation and information processing are incorporated, in analogy of a run-and-tumble bacterium equipped with a ‘brain’. After deriving the generalised fluctuation theorem setting the relation between the entropy production and the information flow in the system, we discuss the optimal feedback control strategies regarding the first passage steps and the total energy consumption of the system. The performed analyses can be directly generalised to other systems with different optimisation goals such as minimal excess work [62, 63] and dissipated heat [63, 64]. We anticipate the concepts developed here to be applied for uncovering the physical rules underlying more complicated responses of regulatory living systems including collective dynamics [65, 66, 67, 68, 69, 70, 71], and also to be useful in future industrial designs of smart active systems with desired physical responses.
Acknowledgements.
F.M. acknowledges the supports by the National Natural Science Foundation of China (Grant No. 12275332 and 12447101), Wenzhou Institute (Grant No. WIUCASQD2023009), and Beijing National Laboratory for Condensed Matter Physics (Grant No. 2023BNLCMPKF005).
     Appendix
The probability, $p(n^{\prime}_{k})$. In the table, $p(n^{\prime}_{k})$, utilised for analysing the mean first passage steps and the mean total energy consumption, depends on both the magnetic field and the measurement error explicitly. In the table, $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L,R)=1/2$ denotes the equilibrium distribution of $n_{k}$ and $n_{k}^{\prime}$ without applying magnetic field, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$, respectively, without applying magnetic field. $p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=L)=\exp[-mB/(k_{B}T)]/\{\exp[-mB/(k_{B}T)]+\exp[mB/(k_{B}T)]\}$ and $p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=R)=1-p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=L)$ denote the equilibrium distribution of $n_{k}^{\prime}$ with applied magnetic field, and the transition rate is obtained as in the Kramers process [72, 73], $k_{M}=k_{M}^{L\rightarrow R}+k_{M}^{R\rightarrow L}=k_{0}\cosh[mB/(k_{B}T)]$ with applied magnetic field.
$n^{\prime}_{k}$ $p(n^{\prime}_{k})$
$L$ $p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)+p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)\epsilon$
$p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)-p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)(1-\epsilon)+[p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)+p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)(1-\epsilon)$
$p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)\epsilon$
$R$ $p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)-p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)\epsilon$
$p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)+p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)(1-\epsilon)+[p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)(1-\epsilon)$
$p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)+p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)\epsilon$
Table 1: The probability, $p(n^{\prime}_{k})$.
The probability, $p(n^{\prime}_{k}|n_{k},m_{k})$. Here we show an important probability utilised in the main text, $p(n^{\prime}_{k}|n_{k},m_{k})$. For other probabilities, one can refer to supplemental materials for their explicit expressions. [53].
$n_{k},m_{k},n^{\prime}_{k}$ $p(n^{\prime}_{k}|n_{k},m_{k})$
$L,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$
$L,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$
$R,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$
$R,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$
$L,L,L$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)+[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)]\cdot e^{-k_{M}\Delta t}$
$L,L,R$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)-[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)]\cdot e^{-k_{M}\Delta t}$
$R,L,L$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)-[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)]\cdot e^{-k_{M}\Delta t}$
$R,L,R$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)+[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)]\cdot e^{-k_{M}\Delta t}$
Table 2: The probability, $p(n^{\prime}_{k}|n_{k},m_{k})$ in the regulatory process.